DISPERSED FLUORESCENCE SPECTROSCOPY OF THE B̃ ^{2}E′ – X̃ ^{2}A′_{2} TRANSITION OF JET COOLED 14NO_{3} and ^{15}NO_{3}

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated NO_{3} in supersonic free jet expansions and observed laser induced fluorescence ( LIF ) of the B̃ ^{2}E′ – X̃ ^{2}A′_{2} transition. We have measured LIF excitation spectra and dispersed fluorescence ( DF ) spectra from the single vibronic levels ( SVL’s ) of the B̃ ^{2}E′ state of ^{14}NO_{3} and ^{15}NO_{3}. The vibrational structure of the X̃ ^{2}A′_{2} state has been analyzed by comparing the vibrational structures of the DF spectra of the two isotopomers. The 1,053 cm^{−1} band of 14NO_{3} is observed as two bands at 1,039 and 1,053 cm^{−1} with an intensity ratio of 4 : 5, respectively, for ^{15}NO_{3}, which are observed in the DF spectra with our standard resolution ( ∼ 7 cm^{−1} in FWHM ). Higher resolution measurements ( ∼ 2 cm^{−1} in FWHM ) of the DF spectra show that the 1,053 cm^{−1} band of ^{14}NO_{3} is also observed as two bands at 1,051 and 1,056 cm^{−1} with an intensity ratio of 5 : 3, respectively. The 1,051 cm^{−1} band is attributed to be the ν_{1} ( a_{1}’ ) fundamental, because of its little isotope shift. There are two possibilities for another band, the band at 1,056 and 1,038 cm^{−1} for ^{14}NO_{3} and ^{15}NO_{3}, respectively; (1) the ν_{3} ( e′ ) fundamental band^{a}, and (2) the ν_{2} + ν_{4} ( a′′_{2} and e′, respectively ) combination band. If this is the case (1), the ν_{3} band should be observed in IR spectrum, but it has yet to be observed. If (2), the intensity must be stolen from the B̃ ^{2}E′ – Ã ^{2}E′′ transition through the ν_{2} mode, the considerable transition moment of which has been predicted^{b}. A simple consideration for the vibronic coupling^{c} between the Ã ^{2}E′′ and X̃ ^{2}A′_{2} states through the ν_{2} mode can understand about 20 % of the combination band intensity to that of the ν_{1} fundamental. The higher resolution measurements of the DF spectra also show that the 1,499 cm^{−1} band of 14NO_{3} is much stronger than the 1,492 cm^{−1} band in the electronic spectrum, while the latter is the strongest band in the IR absorption spectrum^{d}.

aJ. F. Stanton, J. Chem. Phys. 126, 134309 (2007).

bJ. F. Stanton and M. Okumura, Phys. Chem. Chem. Phys. 11, 4742 (2009).

cE. Hirota, K. Kawaguchi, T. Ishiwata, and I. Tanaka, J. Chem. Phys. 95, 771 (1991).

dT. Ishiwata, I. Tanaka, K. Kawaguchi, and E. Hirota, J. Chem. Phys. 82, 2196 (1985).